Find the vertex, line of symmetry to this problem:
f(x)= 2x^2 - 12x + 11

Answer by stanbon(75887)

Find the vertex, line of symmetry to this problem:
f(x)= 2x^2 - 12x + 11
----
Complete the square:
y = 2(x^2-6x) + 11
---
y = 2(x^2-6x+9) + 11 - 18
---
y = 2(x+3)^2-7
---
Vertex: (-3,-7)
Line of symmetry: x = -3
---------------------------------
Cheers,
Stan H.